### v1

@@ Appendix A Related Work
Both machine learning techniques require crafting high-quality features that capture the important characteristics of programs, a process known as feature engineering. The most prevalent feature vectors are based on the frequencies of various types of instructions within the programs magni2014automatic ; huang2019autophase , designed by expert intuitions. Numerous studies have aimed to reduce the cost of feature design. Following the success of word2vec embeddings in natural language processing mikolov2013distributed , methods like code2vec alon2019code2vec , inst2vec ben2018neural , and IR2vec venkatakeerthy2020ir2vec represent programs as distributed vectors that capture syntactic and semantic information from the abstract syntax tree (AST) or intermediate representation (IR). The surge in deep learning has enabled feeding raw information such as AST alon2018general , control-data flow graphs (CFG) brauckmann2020compiler ; cummins2020programl , and code token sequences cummins2017end into powerful deep neural networks, capable of learning useful representations end-to-end.
@@
Platforms that expose the compiler as a playground for AI experiments have significantly reduced the entry barriers to intelligent compiler research. OpenTuner ansel2014opentuner and YaCoS filho2018yet serve as autotuning frameworks with a range of compiler optimization techniques. Our experiments utilize CompilerGym cummins2022compilergym , which offers user-friendly interfaces for researchers to interact with compilers in a reinforcement learning manner. We are optimistic that the future release of our trained code optimization agents, in conjunction with these platforms, can have a democratizing effect on applying AI techniques to compiler optimizations.
@@ B.1 Compiler Environment
Our experiments are conducted on the CompilerGym platform cummins2022compilergym , version 0.2.5, with LLVM-10.0.0 integration.
@@ B.2 Features and Actions
Table 2: Descriptions of 56-dimension Autophase features (adapted from cummins2022compilergym ).
@@ 54 ArgsPhi Total arguments to Phi nodes
We extract program features following the approach of Autophase huang2019autophase . As described in Section 5.1, we use a 56-dimension Autophase feature vector concatenated with a 42-dimension action histogram vector as the observation. The Autophase feature contains various statistics of the IR code, with each dimension thoroughly explained in Table 2. The action histogram vector contains the counts of actions already taken by the agent within the current episode. Both vectors are normalized to ensure that their values fall within a reasonable range. Specifically, each element of the Autophase vector is divided by the total instruction count of the program, whereas the action histogram vector is normalized by the total action count per episode, set as 45.
Our action space is derived from Autophase huang2019autophase , comprising 45 LLVM optimization passes. However, CompilerGym excludes 3 actions due to updates in the latest LLVM version. Thus, we utilize a total of 42 actions. Table 3 presents the names of the LLVM optimization passes corresponding to these 42 actions.
Table 3: A list of LLVM transformation passes selected as actions.
@@
B.3 Hyperparameters
The hyperparameters for our DreamerV3 agent implementation are outlined in Table 4. For hyperparameters not specified, we use the same value as the original DreamerV3 hafner2023mastering .
Table 4: Hyperparameters of DreamerV3 in our experiments.
@@ B.3 Hyperparameters
Reward smoothing $\alpha$ lee2023dreamsmooth 0.6
@@
B.4 Model-Free Baselines
We use RLlib liang2018rllib to train and test model-free reinforcement learning algorithms including PPO schulman2017proximal , A2C mnih2016asynchronous , IMPALA espeholt2018impala , APEX horgan2018distributed , and DQN mnih2015human . All the experiments are conducted with 10 CPUs and an RTX-3090 GPU over a minimum duration of 10 hours. The training process uses 5 rollout workers to interact with the environment and use 4 evaluation workers to evaluate the checkpoint on the validation set. We use default hyperparameters of algorithms in RLlib following the CompilerGym platform cummins2022compilergym , except that we have carefully tuned the hyperparameters for our strongest baseline, PPO, as listed in Table 5.
Table 5: Hyperparameters for the PPO baseline, well tuned on our dataset to be deviating from the default value in RLlib.
@@ Appendix C Benchmarks
To further test the generalization ability of our CodeZero agent on different programming languages, we borrow the method from FormAI tihanyi2023formai and generate a dataset containing 50 unique Objective-C programs using GPT-3.5. We use the same prompt as FormAI, except that we add an instruction to ask GPT to generate programs that can be directly compiled under Clang version 10.0.0 and do not use ARC (Automatic Reference Counting), to improve the compilation pass rate of generated programs. We compile the generated programs using Clang without including any third-party libraries, and all programs that cannot pass compilation are discarded.
@@ D.1 Learning Curves
The zero-shot test performance of our CodeZero agents during training is shown in Figure 6. Note that we validate and test the agent every 10000 environment steps and report the test performance from the checkpoint that achieved the best validation results for comparison among various methods.
Figure 6: Zero-shot test performance of the CodeZero agent during training. We report mean and standard deviation across three runs.
@@ D.2 Quantitative Results
BLAS geomean 0.931 0.821$\pm$0.018 0.960$\pm$0.004 0.993$\pm$0.005 0.991$\pm$0.000
min 0.707 0.432$\pm$0.106 0.723$\pm$0.076 0.831$\pm$0.148 0.913$\pm$0.000
max 1.000 1.002$\pm$0.004 1.031$\pm$0.010 1.025$\pm$0.005 1.016$\pm$0.000
cBench geomean 0.481 0.737$\pm$0.015 0.858$\pm$0.020 N/A 1.036$\pm$0.002
min 0.299 0.401$\pm$0.097 0.560$\pm$0.030 N/A 0.962$\pm$0.001
max 0.626 1.150$\pm$0.113 1.298$\pm$0.083 N/A 1.395$\pm$0.054
CHStone geomean 0.487 0.701$\pm$0.043 1.037$\pm$0.006 N/A 1.094$\pm$0.005
min 0.402 0.278$\pm$0.108 0.889$\pm$0.044 N/A 1.000$\pm$0.006
max 0.655 1.038$\pm$0.033 1.304$\pm$0.038 N/A 1.378$\pm$0.021
Linux geomean 0.988 0.981$\pm$0.004 1.001$\pm$0.000 0.993$\pm$0.000 0.986$\pm$0.005
min 0.615 0.740$\pm$0.126 1.000$\pm$0.000 0.700$\pm$0.001 0.642$\pm$0.041
max 1.011 1.001$\pm$0.002 1.011$\pm$0.000 1.011$\pm$0.000 1.009$\pm$0.003
Mibench geomean 0.389 0.812$\pm$0.022 1.005$\pm$0.001 N/A 1.006$\pm$0.002
min 0.278 0.352$\pm$0.074 0.858$\pm$0.010 N/A 0.879$\pm$0.003
max 0.760 1.392$\pm$0.074 1.603$\pm$0.015 N/A 1.588$\pm$0.000
NPB geomean 0.530 0.833$\pm$0.011 1.074$\pm$0.010 1.075$\pm$0.017 1.108$\pm$0.001
min 0.191 0.414$\pm$0.043 0.805$\pm$0.104 0.810$\pm$0.082 0.886$\pm$0.000
max 1.066 1.848$\pm$0.193 2.141$\pm$0.062 2.315$\pm$0.028 2.343$\pm$0.012
OpenCV geomean 0.981 0.949$\pm$0.011 1.080$\pm$0.001 1.087$\pm$0.007 1.092$\pm$0.000
min 0.833 0.603$\pm$0.133 0.888$\pm$0.021 0.898$\pm$0.001 0.897$\pm$0.000
max 1.370 1.409$\pm$0.167 1.571$\pm$0.022 1.635$\pm$0.186 1.556$\pm$0.000
TensorFlow geomean 0.983 0.912$\pm$0.009 1.006$\pm$0.003 1.032$\pm$0.000 1.032$\pm$0.001
min 0.927 0.618$\pm$0.046 0.877$\pm$0.009 0.970$\pm$0.010 0.968$\pm$0.011
max 1.010 1.235$\pm$0.034 1.267$\pm$0.007 1.289$\pm$0.008 1.282$\pm$0.002
D.3 Data Distribution Visualization
In Figure 7, we visualize the distribution of our training and test datasets. To accurately represent the dynamic behavior of programs, we randomly select 1000 action sequences, each with a length of 45, from our action space. These sequences are subsequently executed on each program, with the resulting Autophase features concatenated to form a feature vector with dimensions of $1000\times 45\times 56$ for every program. These comprehensive feature vectors are finally processed using t-SNE van2008visualizing for dimensionality reduction and visualization.
Figure 7 illustrates that our training data (denoted as triangles) has a broad coverage of test programs (denoted as circles). This contrasts with the Csmith dataset (denoted as crosses) employed in CompilerGym experiments cummins2022compilergym , which shows a significant deviation from real-world applications. Nonetheless, our visualization can still not perfectly capture the transferability across datasets. For instance, empirical evidence suggests that CodeContests are the most effective in generalizing to OpenCV and TensorFlow, while the visual analysis does not directly imply this.
Figure 7: t-SNE van2008visualizing visualization of programs from training and test datasets.

### v2

@@ Appendix A Extended Related Work
Machine learning techniques introduced in Section 2 require crafting high-quality features that capture the important characteristics of programs, a process known as feature engineering. The most prevalent feature vectors are based on the frequencies of various types of instructions or patterns within the programs (Magni, Dubach, and O’Boyle 2014; Haj-Ali et al. 2020), designed by expert intuitions. Numerous studies have aimed to reduce the cost of feature design. Following the success of word2vec embeddings in natural language processing (Mikolov et al. 2013), methods like code2vec (Alon et al. 2019), inst2vec (Ben-Nun, Jakobovits, and Hoefler 2018), and IR2vec (VenkataKeerthy et al. 2020) represent programs as distributed vectors that capture syntactic and semantic information from the abstract syntax tree (AST) or intermediate representation (IR). The surge in deep learning has enabled feeding raw information such as AST (Alon et al. 2018), control-data flow graphs (CFG) (Brauckmann et al. 2020; Cummins et al. 2021), and code token sequences (Cummins et al. 2017) into powerful deep neural networks, capable of learning useful representations end-to-end.
Large Language Model
With the rise of large language models (LLMs), recent work has introduced LLMs into compiler optimization problems, demonstrating remarkable effectiveness in predicting optimal optimization sequences and comprehending the optimization process (Cummins et al. 2023). LLMs can directly analyze the intermediate representation of code in textual form, enabling a better understanding of the code’s intricate details. However, due to limitations on input length, it has to split programs into shorter functions for training and prediction, preventing it from predicting on code of arbitrary length. We also want to emphasize that optimizing individual functions is quite different from optimizing an entire program. Many optimization passes, such as –inline and –mergefunc, which handle function inlining and merging, are either ineffective or less critical when there’s only one function to optimize.
Frameworks
@@
Platforms that expose the compiler as a playground for AI experiments have significantly reduced the entry barriers to intelligent compiler research. OpenTuner (Ansel et al. 2014), and YaCoS (Filho, Rodriguez, and da Silva 2018) serve as autotuning frameworks with a range of compiler optimization techniques. Our experiments utilize CompilerGym (Cummins et al. 2022), which offers user-friendly interfaces for researchers to interact with compilers in a reinforcement learning manner. We are optimistic that the future release of our compiler world model and code optimization agents, in conjunction with these platforms, can have a democratizing effect on applying AI techniques to compiler optimizations.
@@ B.1 Compiler Environment
Our experiments are conducted on the CompilerGym platform (Cummins et al. 2022), version 0.2.5, with LLVM-10.0.0 integration.
@@ 54 ArgsPhi Total arguments to Phi nodes
Table 3: Descriptions of 56-dimension Autophase features (adapted from (Cummins et al. 2022)).
We extract program features following the approach of Autophase (Haj-Ali et al. 2020). We use a 56-dimension Autophase feature vector concatenated with a 42-dimension action histogram vector as the observation. The Autophase feature contains various statistics of the IR code, with each dimension thoroughly explained in Table 3. The action histogram vector contains the counts of actions already taken by the agent within the current episode. Both vectors are normalized to ensure that their values fall within a reasonable range. Specifically, each element of the Autophase vector is divided by the total instruction count of the program, whereas the action histogram vector is normalized by the total action count per episode, set as 45.
We choose to use this feature extraction method because it is both efficient and effective. In our preliminary experiments, we find that more complex feature extraction methods, such as ProGraML (Cummins et al. 2021) and inst2vec (Ben-Nun, Jakobovits, and Hoefler 2018), significantly slow down the training and inference of CompilerDream, with only marginal improvements in performance and generalization.
We employ two distinct action spaces in our experiments to align with the baseline methods. In Sections 4.2 and Section 4.3, we utilize the full action space comprising all 124 optimization passes in LLVM. In Section 4.4 and Section 4.5, CompilerDream and the baseline methods are trained and tested on an action space derived from Autophase (Haj-Ali et al. 2020). The original version of this action space included 45 LLVM optimization passes. However, CompilerGym excludes 3 actions due to updates in the latest LLVM version. Thus, we utilize a total of 42 actions. Table 4 presents the names of the LLVM optimization passes corresponding to these 42 actions. Using this limited action space is a prevailing practice, and it has been proved to be effective in prior works(Haj-Ali et al. 2020; Cummins et al. 2022). It also makes the training easier since the model can focus on the most effective optimization passes.
@@
Table 4: A list of LLVM transformation passes selected in Autophase’s action space.
B.3 Reward Function
The reward function is defined as the normalized change of the optimization metric $C(s)$:
$r_{t+1}=\frac{C(s_{t})-C(s_{t+1})}{C(s_{0})-C(s_{\mathrm{b}})},$ (9)
where a lower $C$ indicates better performance. In line with other works focused on optimizing code size, we define $C(s)$ as the IR instruction count at compiler state $s$. The $C(s_{\mathrm{b}})$ stands for the performance of a baseline policy, specifically the built-in -Oz flag in our experiments. A total reward greater than 1 means that the optimization sequence performs better than the baseline policy.
B.4 Hyperparameters
The hyperparameters for our DreamerV3 agent implementation are outlined in Table 5. For hyperparameters not specified, we use the same value as the original DreamerV3 (Hafner et al. 2023). Most of the listed hyperparameter values are directly taken from DreamerV2 (Hafner et al. 2021) or previous works, with minimal tuning. However, the loss scales are carefully tuned, as the observation loss is relatively small compared to the reward loss in our setting, and the balance between these losses has a significant impact on CompilerDream’s performance. All experiments share the same set of hyperparameters unless otherwise specified.
@@ B.4 Hyperparameters
Reward smoothing $\alpha$ (Lee, Abbeel, and Lee 2024) 0.6
@@
Table 5: Hyperparameters of DreamerV3 in our experiments.
B.5 Training Model-free Baselines
We use RLlib (Liang et al. 2018) to train and test model-free reinforcement learning algorithms including PPO (Schulman et al. 2017) (Haj-Ali et al. 2020), A2C (Mnih et al. 2016), IMPALA (Espeholt et al. 2018), APEX (Horgan et al. 2018), and DQN (Mnih et al. 2015). All the experiments are conducted with 10 CPUs and an RTX-3090 GPU. For the PPO baseline in Section 4.4, we trained it for around 7 hours since we observed overfit to CodeContests when training for a longer time. Other model-free methods are trained for at least 10 hours. The training process uses 5 workers to interact with the environment and use 4 evaluation workers to evaluate the checkpoint on the validation set. We use default hyperparameters of algorithms in RLlib following the CompilerGym platform (Cummins et al. 2022), except that we have carefully tuned the hyperparameters for PPO, as listed in Table 6. We explored roughly 2 or 3 values for each hyperparameter listed and selected the set of hyperparameters that yielded the best performance on the validation set. The results of the PPO baseline in Section 4.5 and Figure 6 are exactly the same as those in Section 4.4 and Figure 3.
@@ B.5 Training Model-free Baselines
Table 6: Hyperparameters for the PPO baseline, well-tuned on our dataset to be deviating from the default value in RLlib.
B.6 Autotuning
In Section 4.2, we extensively trained CompilerDream on the cBench dataset to evaluate its effectiveness as an autotuning method in comparison to other approaches on CompilerGym’s leaderboard. Similar to other studies, we trained CompilerDream using all programs in cBench, excluding the one called ghostscript due to its considerable size, which would significantly hinder the training process. The action space is set to the full set of 124 actions in LLVM. All other training details remained unchanged.
B.7 Guided Search
The guided search strategy we use in Section 4.2 followed the design in the PPO+Guided Search baseline. During the search, actions are sampled from the agent’s learned policy $\pi_{\psi}(a_{t}|s_{t})$ up to 45 steps, at which point a new episode begins. Code size reductions after applying each action are all recorded to calculate the maximum. To enhance exploration, we add a uniform distribution to the policy, implemented by a $5\%$ chance of selecting a uniformly random action instead of sampling from the policy. At each step, the code size reduction achieved so far is recorded for later comparison, and the elapsed time is monitored to ensure that the search is completed within the 1-minute time limit per benchmark. Compared to the guided search in PPO+Guided Search (in Table 1), we use a horizon of 45 steps instead of 200 steps. Additionally, PPO+Guided Search includes an extra search of 500 steps on the best sequence found, which we do not include, resulting in our wall time being a few seconds shorter than theirs.
B.8 Value Estimation
Here, we provide additional details about the experiment in Section 4.3. The baseline method, Coreset-NVP (Liang et al. 2023), trains a value prediction model that scores sequences from a fixed set of optimization sequences, referred to as the core set, which is obtained through an extensive search on their training set. For each program to optimize, it enumerates the first 45 passes in the top-scored sequences and validates their effectiveness through actual compilation. Specifically, sequences are processed in descending order of scores, executing each pass within a sequence. The compiler state (i.e., the IR) resets to its initial state after all passes in a sequence have been enumerated, before moving on to the next sequence. The code size reduction achieved after each pass is recorded, and the pass, along with its corresponding sequence prefix that yields the highest code size reduction, is selected. The core set in Coreset-NVP contains 50 sequences comprising a total of 625 passes, resulting in an average sequence length of 12.5 passes. Thus, roughly 3 or 4 sequences are attempted per benchmark, given the limit of 45 passes. These passes correspond to the optimization actions defined in Section 3.1, with a total of 79 different optimization passes included.
For our experiment, we set the action space to the full set of all 124 optimization passes available in LLVM when training our world model. Focusing on predicting the target value of cumulative reward after executing each optimization pass in the core set, we replace CompilerDream’s optimization agent with a policy that sequentially outputs the sequences from the core set, allowing the world model to focus on the problem dynamics driven by these sequences. This setup is also comparable to the baseline, as Coreset-NVP employs a value predictor that outputs a 50-dimension vector corresponding to the normalized value of each sequence in the core set. That prediction model is obtained through supervised training.
We train our world model using the CodeContests dataset introduced in Section 3.4, consistent with the experiments in Section 4.4. During training, we increase the weight of the reward loss to 100.0, while keeping the other hyperparameters unchanged. To align with the methodology used in Coreset-NVP, we validate the top 45 sequence prefixes of each program with the highest predicted values through actual compilation and select the best-performing sequence. The world model is trained using three different random seeds, and the average results are reported in Table 2.
B.9 Hardware and Training Time
We train all CompilerDream-based methods with 64 CPUs and an RTX-3090 GPU. In Section 4.4, we trained CompilerDream on the CodeContests dataset for around 1 day and 20 hours. In Sections 4.2 and 4.3, CompilerDream was trained for about 1 day.
The random search baseline in Figure 3 is conducted with 4 CPUs, which is sufficient since it is a single-thread program.
The random search baseline in Table 1 utilizes 80 CPUs, as specified in the write-up attached to CompilerGym’s leaderboard (Facebook 2022). Our CompilerDream + Guided Search method in Table 1 is tested on the same machine used for training CompilerDream, equipped with 64 CPUs and an RTX-3090 GPU.
We ensure that sufficient memory is available for all experiments. Memory size should have little impact on the results in all of our experiments.
B.10 Random Seeds
All of our experiments that report a min-max range or standard deviation were replicated using three different random seeds. These seeds were randomly selected, and the results are generally reproducible with other seeds as well.
@@ Appendix C Benchmarks
To further test the generalization ability of our CompilerDream agent on different programming languages, we borrow the method from FormAI (Tihanyi et al. 2023) and generate a dataset containing 50 unique Objective-C programs using GPT-3.5. We use the same prompt as FormAI, except that we add an instruction to ask GPT to generate programs that can be directly compiled under Clang version 10.0.0 and do not use ARC (Automatic Reference Counting), to improve the compilation pass rate of generated programs. We compile the generated programs using Clang without including any third-party libraries, and all programs that cannot pass compilation are discarded.
@@ Appendix D Extended Experimental Results
Dataset Program Pass Sequences by CompilerDream Agent Code size
O0 CompilerDream Oz (Reduction)
cBench sha -sroa -gvn -sroa -sroa -instcombine -simplifycfg -sroa -functionattrs -licm -functionattrs -early-cse -simplifycfg -indvars -gvn -simplifycfg -early-cse -functionattrs -reassociate -memcpyopt -gvn -simplifycfg -reassociate -functionattrs -early-cse -instcombine -early-cse -lowerinvoke -simplifycfg -functionattrs -functionattrs -memcpyopt -gvn -simplifycfg -functionattrs -reassociate -early-cse -instcombine -early-cse -strip -strip -early-cse -strip -early-cse -strip -early-cse 799 349 500 (1.43$\times$)
bzip2 -sroa -gvn -sroa -sroa -simplifycfg -strip -strip -instcombine -simplifycfg -early-cse -simplifycfg -strip -strip -reassociate -reassociate -memcpyopt -jump-threading -memcpyopt -functionattrs -memcpyopt -functionattrs -licm -functionattrs -gvn -simplifycfg -functionattrs -functionattrs -reassociate -memcpyopt -early-cse -jump-threading -memcpyopt -early-cse -strip -strip -instcombine -simplifycfg -jump-threading -jump-threading -early-cse -jump-threading -early-cse -simplifycfg -strip -functionattrs 28748 13565 15946 (1.18$\times$)
OpenCV #41 -sroa -early-cse -early-cse -simplifycfg -early-cse -early-cse -early-cse -sink -sink -sink -loop-deletion -loop-deletion -loop-deletion -lowerinvoke -simplifycfg -sroa -sroa -sroa -early-cse -early-cse -early-cse -early-cse -lowerinvoke -lowerinvoke -lowerinvoke -globalopt -sroa -instcombine -simplifycfg -early-cse -early-cse -early-cse -sink -globalopt -sroa -sroa -functionattrs -simplifycfg -early-cse -functionattrs -early-cse -sink -functionattrs -functionattrs -functionattrs 28 18 28 (1.56$\times$)
#9 -sroa -gvn -sroa -early-cse -simplifycfg -strip -strip -loop-deletion -lowerinvoke -simplifycfg -sroa -early-cse -instcombine -simplifycfg -early-cse -memcpyopt -early-cse -strip -strip -strip -functionattrs -instcombine -simplifycfg -jump-threading -simplifycfg -memcpyopt -early-cse -simplifycfg -reassociate -memcpyopt -functionattrs -functionattrs -reassociate -memcpyopt -early-cse -instcombine -simplifycfg -jump-threading -early-cse -jump-threading -strip -simplifycfg -early-cse -simplifycfg -functionattrs 9510 6341 9269 (1.18$\times$)
TensorFlow #17 -sroa -gvn -early-cse -simplifycfg -early-cse -early-cse -instcombine -early-cse -sink -sink -sink -lowerinvoke -simplifycfg -sroa -early-cse -memcpyopt -early-cse -memcpyopt -instcombine -simplifycfg -early-cse -early-cse -early-cse -simplifycfg -early-cse -early-cse -strip -strip -indvars -simplifycfg -early-cse -memcpyopt -early-cse -instcombine -early-cse -early-cse -simplifycfg -early-cse -memcpyopt -early-cse -early-cse -instcombine -simplifycfg -early-cse -early-cse 5512 4247 5450 (1.28$\times$)
Table 8: Top performances of the zero-shot CompilerDream agent on individual programs. Full IRs of these programs before and after optimizations are shown in the Supplementary Information.
@@ D.1 Learning Curves
The zero-shot test performance of our CompilerDream agents during training is shown in Figure 7. Note that we validate and test the agent every 10000 environment steps and report the test performance from the checkpoint that achieved the best validation results for comparison among various methods.
Figure 7: Zero-shot test performance of the CompilerDream agent during training. We report mean and standard deviation across three runs.
@@ D.2 Quantitative Results
D.3 Program Case Study
In Addition to the predicted optimization trajectory showcase in our paper, we present more example of optimization results here. Table 8 presents the agent’s top performance outcomes on various benchmark datasets. We observe that the agent indeed produces a specialized optimization strategy tailored for each program. Additionally, these results also highlight certain passes, such as -sroa (scalar replacement of aggregates), -early-cse(early common subexpression elimination) and -simplifycfg, as particularly effective in code size optimization.
Figure 8: t-SNE (Van der Maaten and Hinton 2008) visualization of programs from training and test datasets.
D.4 Data Distribution Visualization
In Figure 8, we visualize the distribution of our training and test datasets. To accurately represent the dynamic behavior of programs, we randomly select 1000 action sequences, each with a length of 45, from our action space. These sequences are subsequently executed on each program, with the resulting Autophase features concatenated to form a feature vector with dimensions of $1000\times 45\times 56$ for every program. These comprehensive feature vectors are finally processed using t-SNE (Van der Maaten and Hinton 2008) for dimensionality reduction and visualization.
Figure 8 illustrates that our training data (denoted as triangles) has a broad coverage of test programs (denoted as circles). This contrasts with the Csmith dataset (denoted as crosses) employed in CompilerGym experiments (Cummins et al. 2022), which shows a significant deviation from real-world applications. Nonetheless, our visualization can still not perfectly capture the transferability across datasets. For instance, empirical evidence suggests that CodeContests are the most effective in generalizing to OpenCV and TensorFlow, while the visual analysis does not directly imply this.
@@ D.4 Data Distribution Visualization
BLAS geomean 0.931 0.821$\pm$0.018 0.960$\pm$0.004 0.993$\pm$0.002 0.993$\pm$0.005 0.991$\pm$0.000
min 0.707 0.432$\pm$0.106 0.723$\pm$0.076 0.916$\pm$0.004 0.831$\pm$0.148 0.913$\pm$0.000
max 1.000 1.002$\pm$0.004 1.031$\pm$0.010 1.014$\pm$0.001 1.025$\pm$0.005 1.016$\pm$0.000
cBench geomean 0.481 0.737$\pm$0.015 0.858$\pm$0.020 1.010$\pm$0.004 N/A 1.036$\pm$0.002
min 0.299 0.401$\pm$0.097 0.560$\pm$0.030 0.817$\pm$0.109 N/A 0.962$\pm$0.001
max 0.626 1.150$\pm$0.113 1.298$\pm$0.083 1.405$\pm$0.014 N/A 1.395$\pm$0.054
CHStone geomean 0.487 0.701$\pm$0.043 1.037$\pm$0.006 1.071$\pm$0.002 N/A 1.094$\pm$0.005
min 0.402 0.278$\pm$0.108 0.889$\pm$0.044 0.967$\pm$0.027 N/A 1.000$\pm$0.006
max 0.655 1.038$\pm$0.033 1.304$\pm$0.038 1.313$\pm$0.045 N/A 1.378$\pm$0.021
Linux geomean 0.988 0.981$\pm$0.004 1.001$\pm$0.000 0.985$\pm$0.001 0.993$\pm$0.000 0.986$\pm$0.005
min 0.615 0.740$\pm$0.126 1.000$\pm$0.000 0.615$\pm$0.000 0.700$\pm$0.001 0.642$\pm$0.041
max 1.011 1.001$\pm$0.002 1.011$\pm$0.000 1.009$\pm$0.003 1.011$\pm$0.000 1.009$\pm$0.003
Mibench geomean 0.389 0.812$\pm$0.022 1.005$\pm$0.001 1.000$\pm$0.007 N/A 1.006$\pm$0.002
min 0.278 0.352$\pm$0.074 0.858$\pm$0.010 0.833$\pm$0.062 N/A 0.879$\pm$0.003
max 0.760 1.392$\pm$0.074 1.603$\pm$0.015 1.524$\pm$0.045 N/A 1.588$\pm$0.000
NPB geomean 0.530 0.833$\pm$0.011 1.074$\pm$0.010 1.071$\pm$0.018 1.075$\pm$0.017 1.108$\pm$0.001
min 0.191 0.414$\pm$0.043 0.805$\pm$0.104 0.801$\pm$0.070 0.810$\pm$0.082 0.886$\pm$0.000
max 1.066 1.848$\pm$0.193 2.141$\pm$0.062 2.235$\pm$0.101 2.315$\pm$0.028 2.343$\pm$0.012
OpenCV geomean 0.981 0.949$\pm$0.011 1.080$\pm$0.001 1.082$\pm$0.004 1.087$\pm$0.007 1.092$\pm$0.000
min 0.833 0.603$\pm$0.133 0.888$\pm$0.021 0.897$\pm$0.002 0.898$\pm$0.001 0.897$\pm$0.000
max 1.370 1.409$\pm$0.167 1.571$\pm$0.022 1.620$\pm$0.193 1.635$\pm$0.186 1.556$\pm$0.000
TensorFlow geomean 0.983 0.912$\pm$0.009 1.006$\pm$0.003 1.010$\pm$0.019 1.032$\pm$0.000 1.032$\pm$0.001
min 0.927 0.618$\pm$0.046 0.877$\pm$0.009 0.841±0plus-or-minus0$\pm$ 0± 0.134 0.970$\pm$0.010 0.968$\pm$0.011
max 1.010 1.235$\pm$0.034 1.267$\pm$0.007 1.281$\pm$0.006 1.289$\pm$0.008 1.282$\pm$0.002